OpenStudy (anonymous):

Factor out the greatest common factor from the polynomial 3x+27

OpenStudy (anonymous):

Factor out 3.

OpenStudy (anonymous):

3(x+9)

OpenStudy (anonymous):

so 3(x+9)

OpenStudy (anonymous):

i don't understand

OpenStudy (anonymous):

each term in 3x+27 can be factored as: 3x = 3 * x 27 = 3 * 9 notice that both have a common factor of 3.... and that's what is "factored" out... so 3x+27 = 3(x+9)

OpenStudy (anonymous):

oh ok thank you guys :)